OpenStudy (j2lie):

How do you solve this problem 12x2-22x-20

OpenStudy (anonymous):

Restate the question ? Is it 12 times 2 - 22x or 12x times 2

OpenStudy (j2lie):

12x with the exponent 2.

OpenStudy (anonymous):

Alright :) Next time you type that type it like 12x ^2

OpenStudy (j2lie):

I didn't know I was suppose to do that.

OpenStudy (anonymous):

Just a hint. I don't know how to solve this though sorry :(

OpenStudy (anonymous):

first take common factor 2(6x^2-11x-10) 2(3x+2)(2x-5)

OpenStudy (j2lie):

it's 2(3x+4)(x-15)

OpenStudy (anonymous):

no

OpenStudy (j2lie):

yes because you have to multiply the 6 to the 10.

OpenStudy (anonymous):

no

OpenStudy (j2lie):

if you add -5 to 3 it would be -3 not -11.

OpenStudy (anonymous):

no if u have for example x^2-2x-6=(x-4)(x+2) when u add (-4)(x)+2x=-2x which is the middle term

OpenStudy (anonymous):

use quadratic formula or factor it out

OpenStudy (j2lie):

I factor it with 2.

OpenStudy (anonymous):

yes quadratuc formula @sierraabel thank u very much i just didn't know the vocabulary

OpenStudy (j2lie):

I'm only in the eighth grade.

OpenStudy (anonymous):

what do u call it in english

OpenStudy (j2lie):

factoring.

OpenStudy (j2lie):

what is quadratic formula?

OpenStudy (anonymous):